OpenStudy (anonymous):

What is the slope of the line 8x + 3y = 8 – 2x?

OpenStudy (jhannybean):

From the last problem we did, can you solve for y by subtracting 8x from both sides?

OpenStudy (anonymous):

i have 3y=8-10x now what lol

OpenStudy (jhannybean):

now divide by 3 the Left hand side and LHS of the equation to solve for y

OpenStudy (jhannybean):

left hand side /right hand side*

OpenStudy (anonymous):

okay well im wrong y= 2.6 - 3.3

OpenStudy (jhannybean):

\[\frac{3y}{3}=\frac{8-2x}{3}\]

OpenStudy (anonymous):

wouldnt 8-2x/3 stay the same?

OpenStudy (jhannybean):

\[\frac{\cancel3y}{\cancel3}=\frac{8-2x}{3}\]\[\large y=\frac{8}{3}-\frac{\color{red}2x}{\color{red}3}\]

OpenStudy (anonymous):

for 8/3 i have 2.6 repeating

OpenStudy (jhannybean):

that's your slope :) Your slope is the value that multiplies to your value of x it follows the form \[\large y=\color{red}mx +b\] and in this case your m=2/3

OpenStudy (anonymous):

for 2/3 i have 0.66 repeating.

OpenStudy (jhannybean):

You don't need to put it in decimal form, necessarily

OpenStudy (jhannybean):

You could leave it as a fraction.

OpenStudy (anonymous):

so y=2/3x + 2.6

OpenStudy (jhannybean):

and the 2.6 can stay as 8/3 too.

OpenStudy (anonymous):

y=2/3x+8/3 so this is the answer all together?

OpenStudy (jhannybean):

you're only looking for your slope, right? i mentioned a few posts above where you can find your slope. your slope is multiplied by your "x"

OpenStudy (anonymous):

yes

OpenStudy (jhannybean):

so your slope would be \[\huge y= \color{red}{\frac{2}{3}}x+\frac{8}{3}\] the RED portion,is yoru slope.it is multiplied by your x-value.

OpenStudy (anonymous):

great! it will take a few more of these problems for me to fully get it hahah, thank you very much !

OpenStudy (jhannybean):

no problem, lol.